OpenStudy (anonymous):

Can someone help me figure out this logarithmic equation??? log8(x + 3) = 1 the 8 is tiny

OpenStudy (trojanpoem):

log base 8 for (x+3) = 1 x+3 = 8^1 x = 5

OpenStudy (anonymous):

D: Omg, thanks so much

OpenStudy (trojanpoem):

you're welcome.